# for Large Language Models

### v3

@@ 1 Introduction
Substantial efforts have been dedicated to the task of construction instruction data through human annotations (Wang et al., 2022; Köpf et al., 2023), requiring a significant allocation of resources. Against this backdrop, LLMs are utilized to synthesize large-scale instruction data automatically  (Wang et al., 2023b; Xu et al., 2023; Li et al., 2023). These methods could scale up the size of instruction-following data, but they still inevitably suffer limited diversity and complexity, resulting in an unbalanced distribution and poor quality of instruction data. Recent studies (Zhou et al., 2023; Chen et al., 2023; Xu et al., 2023) have unveiled a seminal revelation, indicating that even a small quantity of high-quality instruction data has the potential to yield robust performance. In general, instruction processing is an important process requiring careful attention to detail and rigorous quality assurance procedures to construct a high-quality instruction dataset for LLMs.
Unfortunately, the availability of open-source tools for instruction processing remains limited, especially in comparison to many open-source projects on models and training infrastructures (Touvron et al., 2023a, b; Taori et al., 2023; Scao et al., 2022; Chiang et al., 2023; Zeng et al., 2023). Existing projects are often high-customized to their own needs, lacking a systematized and modular processing ability to address diverse processing pipelines for LLMs. For instance, the Alpaca (Taori et al., 2023) dataset targets the augmentation of diversity for LLaMA tuning, whereas AlpaGasus (Chen et al., 2023) focuses on filtering out low-quality instances from Alpaca. Thorough development of instruction processing systems for the ever-evolving and emerging requirements of LLM remains unexplored, particularly in light of the quick expansion of inventive LLM applications spanning various fields.
To address this issue, we develop EasyInstruct as depicted in Figure 1, an easy-to-use instruction processing framework for LLMs. Given some existing chat data, corpus, or Knowledge Graphs, EasyInstruct can handle instruction generation, selection, and prompting processes, while also considering their combination and interaction. These consistencies facilitate further development and comparisons of various methods, thus promoting the advancement of better instruction processing work. We further conduct experiments with EasyInstruct to validate its effectiveness in instruction processing. Currently, EasyInstuct is open-sourced on GitHub and has already received around 300 stars. We are committed to the long-term maintenance of EasyInstruct, providing continuous support for new features to ensure its effectiveness as a framework for instruction processing and synthetic data generation (Bauer et al., 2024).
@@ 2 Background
Despite the effectiveness of instruction tuning, constructing high-quality large-scale instructions which effectively encompass the target behaviors remains a non-trivial challenge in this realm. Existing instruction datasets are often limited in terms of diversity, quantity, and creativity, which underscores the significance of instruction processing. One typical method for constructing instruction datasets is data integration. In this method, instructional datasets are constructed by merging existing annotated datasets with descriptions of tasks in natural language (Longpre et al., 2023; Sanh et al., 2022; Anand et al., 2023). Another prevalent method for constructing instruction datasets is automated generation. To alleviate the need for extensive human annotation or manual data gathering, automated methods have been proposed to generate large volumes of instructional data through the use of LLMs. Instructions can be sourced from chat data (Chiang et al., 2023) or expanded on a small set of seed instructions using LLMs (Wang et al., 2023b; Xu et al., 2023; Li et al., 2023). Subsequently, the collected instructions are fed into LLMs to generate corresponding inputs and outputs. In EasyInstruct, our primary focus lies on automated approaches for instruction generation due to their high efficiency and scalability.
Another promising research direction of instruction processing is the selection of high-quality instruction. Recently, numerous studies (Zhou et al., 2023; Chen et al., 2023; Xu et al., 2023; Liu et al., 2023) have investigated the issue of the scale of the instruction dataset for fine-tuning and have indicated that merely increasing the number of instructions may not necessarily result in enhancements. Instead, a modest volume of high-quality instruction data can influence the fine-tuning of LLMs, yielding solid performance. Thus, optimizing the instruction dataset and enhancing its quality play a critical role in fine-tuning LLMs effectively.
@@ 3.2 APIs & Engines
Generators Self-Instruct Chat
The method that randomly samples a few instructions as demonstrations and generates more instructions and input-output pairs using LLM (Wang et al., 2023b).
Evol-Instruct Chat
The method that incrementally upgrades an initial set of instructions into more complex instructions by prompting an LLM with specific prompts (Xu et al., 2023).
Backtranslation Corpus
The method that creates a training instance by predicting an instruction that would be correctly answered by a paragraph in the corpus (Li et al., 2023).
KG2Instruct KG
The method that generates Information Extraction (IE) instruction datasets incorporating existing Knowledge Graphs (Gui et al., 2023).
@@
Selectors Deduplication Statistics-based
Repetitive input and output of instances.
Length Statistics-based
The bounded length of every pair of instruction and output.
MTLD Statistics-based
A metric for assessing the lexical diversity in text, defined as the average length of word sequences that sustain a minimum threshold TTR score (McCarthy and Jarvis, 2010).
ROUGE N-gram-based
Recall-oriented understudy for gisting evaluation (Lin, 2004).
CIRS Structure-based
The score using the abstract syntax tree to encode structural and logical attributes, to evaluate the correlation between code and reasoning abilities (Bi et al., 2023).
Perplexity LM-based
The exponentiated average negative log-likelihood of text.
GPT Score LM-based
The score that ChatGPT/GPT4 assigns to assess how effectively the AI Assistant’s response aligns with the user’s instructions.
@@ 3.3 Generators
Given an unannotated corpus, Instruction Backtranslation (Li et al., 2023) creates an instruction following training instance by predicting an instruction that would be correctly answered by a paragraph in the document or corpus. Considering the mixed quality of human-written web text and the presence of noise in generated content, only the highest quality instances are reserved.
@@
from easyinstruct import SelfInstructGenerator
from easyinstruct import GPTScoreSelector
from easyinstruct.utils.api import set_openai_key
@@
set_openai_key("YOUR-KEY")
[…]
generator = SelfInstructGenerator(
data_format = "alpaca",
seed_tasks_path = "seed_tasks.jsonl",
generated_instances_path = "generation.jsonl",
num_instructions_to_generate=100,
engine = "gpt-3.5-turbo",
[…]
generator.generate()
[…]
selector = GPTScoreSelector(
%**** design.tex Line 150 ****source_file_path = "generation.jsonl",
engine = "gpt-3.5-turbo",
threshold = 4,
[…]
selector.process()
@@ 3.4 Selectors
The Selectors module is designed to streamline the process of choosing instructions, enabling the curation of instruction datasets from raw instruction data. This raw data might originate from openly accessible instruction datasets or be created internally within the framework. Table 1 provides a comprehensive overview of various metrics for instruction quality evaluation. We divide the evaluation metrics into four categories based on the principle of their implementation: statistics-based, n-gram-based, structure-based and LM-based. All Selector classes derive from a common base class, BaseSelector. It includes fundamental attributes and abstract methods such as loading, processing, and dumping of data. In EasyInstruct, multiple Selectors can be grouped for convenient usage, which allows users to achieve more concise and readable code. A running example of using a Selector class is shown in Figure 2.
@@ 3.5 Prompts
The Prompts module standardizes the instruction prompting step, in which user requests are constructed as instruction prompts and sent to specific LLMs to obtain responses. Similar to Selectors, all Prompts classes inherit from a common base class, BasePrompt, which includes necessary attributes and abstract methods. In the mentioned base class, there are functionalities provided for building prompts, requesting generation results from LLMs, and parsing the responses received from LLMs. The base class also provides mechanisms to handle error conditions and exceptions that may occur during the whole process. Users can inherit from the base class and customize or extend its functionality based on their specific requirements. We also equip EasyInstruct with various prompting techniques and application adaptions (e.g. Chain-of-Thought, Information Extraction, Multimodal, etc.) by providing a consistent and standardized interface, enabling efficient instruction prompting for LLMs.
@@ 4 Evaluation
[Output] def calculateCost(unitPrice, quantity, shippingCost, taxRate):
totalPrice = (unitPrice * quantity) + shippingCost
totalTax = totalPrice * taxRate
totalCost = totalPrice + totalTax
return totalCost
@@ 4.1 Experiment Setups
We adopt the popular Self-Instruct (Wang et al., 2023b) and Evol-Instruct (Xu et al., 2023) methods implemented in EasyInstruct to synthetic instruction dataset, containing instructions paired with instance inputs and outputs separately. We mainly consider four instruction datasets as follows: (a) self_instruct_5k is constructed by employing the Self-Instruct method to distill instruction data from text-davinci-003; (b) alpaca_data_5k is randomly sampled from the Alpaca dataset (Taori et al., 2023); (c) evol_instruct_5k is constructed by employing the Evol-Instruct method; (d) easyinstruct_5k is collected by integrating the three instruction datasets above and applying multiple Selectors in EasyInstruct to extract high-quality instruction datasets. We detail the hyper-parameters for selection in a configuration file, as shown in Appendix A.2.1.
@@ 4.2 Experiment Results
To study the diversity of the instruction datasets considered in our experiments, we identify the verb-noun structure in the generated instructions and plot the top 20 most prevalent root verbs and their top 4 direct nouns in Appendix A.4, following the approach of Wang et al. (2023b). Overall, we see a wide range of intents and textual formats within these instructions.
@@
To conduct a qualitative evaluation of EasyInstruct, we sample several instruction examples selected by the Selctors module in easyinstruct_5k for the case study. We also attach the corresponding evaluation scores for each of these instruction examples, as shown in Table 2. We observe that the selected instructions often possess fluent language and meticulous logic.
@@ A.1.1 Installation from GitHub Repository
pip install git+https://github.com/zjunlp/EasyInstruct@main
@@ A.1.2 Installation for Local Development
git clone https://github.com/zjunlp/EasyInstruct
cd EasyInstruct
pip install -e .
@@ A.1.3 Installation from PyPI
pip install easyinstruct
@@ A.2.1 Shell Script
generator: SelfInstructGenerator: target_dir: data/generations/ data_format: alpaca seed_tasks_path: data/seed_tasks.jsonl generated_instructions_path: generated_instructions.jsonl generated_instances_path: generated_instances.jsonl num_instructions_to_generate: 100 engine: gpt-3.5-turbo num_prompt_instructions: 8
@@
selector:
source_file_path:
target_dir: data/selections/
target_file_name: case.jsonl
LengthSelector:
min_instruction_length: 3
max_instruction_length: 150
min_response_length: 1
max_response_length: 350
Deduplicator:
RougeSelector:
threshold: 0.7
%**** appendix.tex Line 125 ****GPTScoreSelector:
engine: gpt-3.5-turbo
threshold: 4
MTLDSelector:
ttr_threshold: 0.72
min_mtld: 8
max_mtld: 22
PPLSelector:
threshold: 200
model_name: gpt2
device: cuda
RandomSelector:
num_instructions_to_sample: 100
seed: 42
@@ A.3 Detailed Hyper-Parameters
A.4 Diversity of the Instruction Datasets
Following Wang et al. (2023b), we illustrate the twenty most common root verbs along with their four primary direct noun associations in the generated instruction datasets considered in our experiments in Figure 10. Overall, we see quite diverse intents and textual formats in these instructions.
A.5 Prompt Template for Instruction Tuning
@@
### Instruction:
{instruction}
### Input:
{input}
@@
A.6 API Services Available in EasyInstruct

### v4

@@ 1 Introduction
Substantial efforts have been dedicated to the task of construction instruction data through human annotations (Wang et al., 2022; Köpf et al., 2023), requiring a significant allocation of resources. Against this backdrop, LLMs are utilized to synthesize large-scale instruction data automatically  (Wang et al., 2023b; Xu et al., 2023; Li et al., 2023b). These methods could scale up the size of instruction-following data, but they still inevitably suffer limited diversity and complexity, resulting in an unbalanced distribution and poor quality of instruction data. Recent studies (Zhou et al., 2023; Chen et al., 2023a; Xu et al., 2023) have unveiled a seminal revelation, indicating that even a small quantity of high-quality instruction data has the potential to yield robust performance. In general, instruction processing is an important process requiring careful attention to detail and rigorous quality assurance procedures to construct a high-quality instruction dataset for LLMs.
Unfortunately, the availability of open-source tools for instruction processing remains limited, especially in comparison to many open-source projects on models and training infrastructures (Touvron et al., 2023a, b; Taori et al., 2023; Scao et al., 2022; Chiang et al., 2023; Zeng et al., 2023). Existing projects are often highly-customized to their own needs, lacking a systematized and modular processing ability to address diverse processing pipelines for LLMs. For instance, the Alpaca (Taori et al., 2023) dataset targets the augmentation of diversity for LLaMA tuning, whereas AlpaGasus (Chen et al., 2023a) focuses on filtering out low-quality instances from Alpaca. Thorough development of instruction processing systems for the ever-evolving and emerging requirements of LLM remains unexplored, particularly in light of the quick expansion of inventive LLM applications spanning various fields.
To address this issue, we develop EasyInstruct as depicted in Figure 1, an easy-to-use instruction processing framework for LLMs. Given some existing chat data, corpus, or Knowledge Graphs, EasyInstruct can handle instruction generation, selection, and prompting processes, while also considering their combination and interaction. These consistencies facilitate further development and comparisons of various methods, thus promoting the advancement of better instruction processing work. We further conduct experiments with EasyInstruct to validate its effectiveness in instruction processing. Currently, EasyInstuct is open-sourced on GitHub and has already received over 300 stars. We are committed to the long-term maintenance of EasyInstruct, providing continuous support for new features to ensure its effectiveness as a framework for instruction processing and synthetic data generation (Bauer et al., 2024).
@@ 2 Background
Despite the effectiveness of instruction tuning, constructing high-quality large-scale instructions which effectively encompass the target behaviors remains a non-trivial challenge in this realm. Existing instruction datasets are often limited in terms of diversity, quantity, and creativity, which underscores the significance of instruction processing. One typical method for constructing instruction datasets is data integration. In this method, instructional datasets are constructed by merging existing annotated datasets with descriptions of tasks in natural language (Longpre et al., 2023; Sanh et al., 2022; Anand et al., 2023). Another prevalent method for constructing instruction datasets is automated generation. To alleviate the need for extensive human annotation or manual data gathering, automated methods have been proposed to generate large volumes of instructional data through the use of LLMs. Instructions can be sourced from chat data (Chiang et al., 2023) or expanded on a small set of seed instructions using LLMs (Wang et al., 2023b; Xu et al., 2023; Li et al., 2023b). Subsequently, the collected instructions are fed into LLMs to generate corresponding inputs and outputs. In EasyInstruct, our primary focus lies on automated approaches for instruction generation due to their high efficiency and scalability.
Another promising research direction of instruction processing is the selection of high-quality instruction. Recently, numerous studies (Zhou et al., 2023; Chen et al., 2023a; Xu et al., 2023; Liu et al., 2023) have investigated the issue of the scale of the instruction dataset for fine-tuning and have indicated that merely increasing the number of instructions may not necessarily result in enhancements. Instead, a modest volume of high-quality instruction data can influence the fine-tuning of LLMs, yielding solid performance. Thus, optimizing the instruction dataset and enhancing its quality play a critical role in fine-tuning LLMs effectively.
@@ 3.2 APIs & Engines
Generators Self-Instruct Chat The method that randomly samples a few instructions as demonstrations and generates more instructions and input-output pairs using LLM (Wang et al., 2023b).
Evol-Instruct Chat The method that incrementally upgrades an initial set of instructions into more complex instructions by prompting an LLM with specific prompts (Xu et al., 2023).
Backtranslation Corpus The method that creates a training instance by predicting an instruction that would be correctly answered by a paragraph in the corpus (Li et al., 2023b).
KG2Instruct KG The method that generates Information Extraction (IE) instruction datasets incorporating existing Knowledge Graphs (Gui et al., 2023).
@@
Selectors Deduplication Statistics-based Repetitive input and output of instances.
Length Statistics-based The bounded length of every pair of instruction and output.
MTLD Statistics-based A metric for assessing the lexical diversity in text, defined as the average length of word sequences that sustain a minimum threshold TTR score (McCarthy and Jarvis, 2010).
ROUGE N-gram-based Recall-oriented understudy for gisting evaluation (Lin, 2004).
CIRS Structure-based The score using the abstract syntax tree to encode structural and logical attributes, to evaluate the correlation between code and reasoning abilities (Bi et al., 2023).
Perplexity LM-based The exponentiated average negative log-likelihood of text.
GPT Score LM-based The score that ChatGPT/GPT4 assigns to assess how effectively the AI Assistant’s response aligns with the user’s instructions.
@@ 3.3 Generators
Given an unannotated corpus, Instruction Backtranslation (Li et al., 2023b) creates an instruction following training instance by predicting an instruction that would be correctly answered by a paragraph in the document or corpus. Considering the mixed quality of human-written web text and the presence of noise in generated content, only the highest quality instances are reserved.
@@
vs {minted}[breaklines, firstnumber=last, bgcolor=bg, fontfamily=courier,fontsize=]python from easyinstruct import SelfInstructGenerator from easyinstruct import GPTScoreSelector from easyinstruct.utils.api import set_openai_key
# Step1: Set your own API-KEY set_openai_key("YOUR-KEY")
# Step2: Declare a generator class generator = SelfInstructGenerator( data_format = "alpaca", seed_tasks_path = "seed_tasks.jsonl", generated_instances_path = "generation.jsonl", num_instructions_to_generate=100, engine = "gpt-3.5-turbo", )
# Step3: Generate self-instruct data generator.generate()
# Step4: Declare a selector class selector = GPTScoreSelector( source_file_path = "generation.jsonl", engine = "gpt-3.5-turbo", threshold = 4, )
# Step5: Process raw data selector.process()
@@ 3.4 Selectors
The Selectors module is designed to streamline the process of filtering instructions, enabling the curation of instruction datasets from raw instruction data. This raw data might originate from publicly accessible instruction datasets or be synthesised in advence by the Generators module. Table 1 provides a comprehensive overview of various metrics for instruction quality evaluation. We divide the evaluation metrics into four categories based on the principle of their implementation: statistics-based, n-gram-based, structure-based and LM-based. All Selector classes derive from a common base class, BaseSelector. It includes fundamental attributes and abstract methods such as loading, processing, and dumping of data. In EasyInstruct, multiple Selectors can be grouped for convenient usage, which allows users to achieve more concise and readable code. A running example of using a Selector class is shown in Figure 2.
@@ 3.5 Prompts
The Prompts module standardizes the instruction prompting step, in which user requests are constructed as instruction prompts and sent to specific LLMs to obtain responses. Utilizing the Prompts module with a series of well-designed and refined prompts enhances the ability of Generators and Selectors to effectively fulfill their respective functions. Similar to Selectors, all Prompts classes inherit from a common base class, BasePrompt, which includes necessary attributes and abstract methods. In the mentioned base class, there are functionalities provided for building prompts, requesting generation results from LLMs, and parsing the responses received from LLMs. The base class also provides mechanisms to handle error conditions and exceptions that may occur during the whole process. Users can inherit from the base class and customize or extend its functionality based on their specific requirements. We also equip EasyInstruct with various prompting techniques and application adaptions (e.g. Chain-of-Thought, Information Extraction, Multimodal, etc.) by providing a consistent and standardized interface, enabling efficient instruction prompting for LLMs.
@@ 4 Evaluation
In terms of evaluation, we will introduce the experiment setups and illustrate the empirical results of multiple modules implemented in EasyInstruct to demonstrate its capability.
@@
[Output] def calculateCost(unitPrice, quantity, shippingCost, taxRate): totalPrice = (unitPrice * quantity) + shippingCost totalTax = totalPrice * taxRate totalCost = totalPrice + totalTax return totalCost
@@ 4.1 Experiment Setups
We adopt the popular Self-Instruct (Wang et al., 2023b) and Evol-Instruct (Xu et al., 2023) methods implemented in EasyInstruct to synthesize instruction datasets, containing instructions paired with instance inputs and outputs separately. We mainly consider four instruction datasets as follows: (a) self_instruct_5k is constructed by employing the Self-Instruct method to distill instruction data from text-davinci-003; (b) alpaca_data_5k is randomly sampled from the Alpaca dataset (Taori et al., 2023); (c) evol_instruct_5k is constructed by employing the Evol-Instruct method; (d) easyinstruct_5k is collected by integrating the three instruction datasets above and applying multiple Selectors in EasyInstruct to extract high-quality instruction datasets. We detail the hyper-parameters for selection in a configuration file, as shown in Appendix A.2.1.
@@ 4.2 Experiment Results
To study the diversity of the instruction datasets considered in our experiments, we identify the verb-noun structure in the generated instructions and plot the top 20 most prevalent root verbs and their top 4 direct nouns in Figure 4, following the approach of Wang et al. (2023b). Overall, we see a wide range of intents and textual formats within these instructions.
@@
To conduct a qualitative evaluation of EasyInstruct, we sample several instruction examples selected by the Selectors module in easyinstruct_5k for the case study. We also attach the corresponding evaluation scores for each of these instruction examples, as shown in Table 2. We observe that the selected instructions often possess fluent language and meticulous logic.
@@ 5 Conclusion and Future Work
Limitations
In this paper, we are committed to unifying all phases of instruction data processing including instruction generation, selection, and prompting. Despite our efforts, this paper may still have some remaining limitations.
The Scope of Instruction Selection Methods.
We implement various instruction selection methods within the Selectors module. Based on the evaluation metrics utilized and the model base employed, the implemented instruction data selection methods can be divided into three categories: methods based on a system of indicators, methods utilizing powerful LLMs like ChatGPT, and methods employing small models (Wang et al., 2024). However, another line of work (Li et al., 2023a, c, b; Wu et al., 2023; Chen et al., 2023b; Kung et al., 2023) employs trainable LLMs like LLaMA for computation formulas in instruction selection processes, which are not integrated into the Selectors module. Although our design choice is to decouple instruction processing and model training into two separate phases, we regard it as a limitation that may be addressed by future work.
Statistics for evaluating efficiency.
In our evaluation, we fine-tune a LLaMA2 (7B) model utilizing multiple modules implemented in EasyInstruct. Compared to models fine-tuned on other instruction datasets constructed without EasyInstruct, our model achieves optimal results, demonstrating EasyInstruct’s capability. Although we also qualitatively demonstrate the ease of writing code for instruction processing with multiple code samples and configuration files using EasyInstruct, a limitation is the lack of appropriate statistics for quantitatively evaluating efficiency.
Acknowledgments
We would like to express gratitude to the anonymous reviewers for their kind comments. This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), Yongjiang Talent Introduction Programme (2021A-156-G), CCF-Baidu Open Fund, Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
@@ A.1.1 Installation from GitHub Repository
[…]
emacs {minted}[breaklines, firstnumber=last, bgcolor=bg, fontfamily=courier,fontsize=]bash pip install git+https://github.com/zjunlp/EasyInstruct@main
@@ A.1.2 Installation for Local Development
[…]
emacs {minted}[breaklines, firstnumber=last, bgcolor=bg, fontfamily=courier,fontsize=]bash git clone https://github.com/zjunlp/EasyInstruct cd EasyInstruct pip install -e .
@@ A.1.3 Installation from PyPI
[…]
emacs {minted}[breaklines, firstnumber=last, bgcolor=bg, fontfamily=courier,fontsize=]bash pip install easyinstruct
@@ A.2.1 Shell Script
[…]
emacs {minted}[breaklines, firstnumber=last, bgcolor=bg, fontfamily=courier,fontsize=]yaml generator: SelfInstructGenerator: target_dir: data/generations/ data_format: alpaca seed_tasks_path: data/seed_tasks.jsonl generated_instructions_path: generated_instructions.jsonl generated_instances_path: generated_instances.jsonl num_instructions_to_generate: 100 engine: gpt-3.5-turbo num_prompt_instructions: 8
@@
[…]
emacs {minted}[breaklines, firstnumber=last, bgcolor=bg, fontfamily=courier,fontsize=]yaml selector: source_file_path: target_dir: data/selections/ target_file_name: case.jsonl LengthSelector: min_instruction_length: 3 max_instruction_length: 150 min_response_length: 1 max_response_length: 350 Deduplicator: RougeSelector: threshold: 0.7 GPTScoreSelector: engine: gpt-3.5-turbo threshold: 4 MTLDSelector: ttr_threshold: 0.72 min_mtld: 8 max_mtld: 22 PPLSelector: threshold: 200 model_name: gpt2 device: cuda RandomSelector: num_instructions_to_sample: 100 seed: 42
@@
emacs {minted}[breaklines, firstnumber=last, bgcolor=bg, fontfamily=courier,fontsize=]bash config_file="" openai_api_key=""
python demo/run.py  –config $config_{f}ile\ --openai_{a}pi_{k}ey$openai_api_key
@@ A.3 Detailed Hyper-Parameters
A.4 Prompt Template for Instruction Tuning
@@
### Instruction: {instruction}
### Input: {input}
@@
A.5 API Services Available in EasyInstruct